Simons Center for Geometry and Physics Arts Program

Report to Board of Trustees

May 14, 2012

1. Administration. Since January 2012 the SCGP Arts Program has been without a Director. Exhibitions have been curated by free-lance personnel, coordinated by the Advisory Committee and administered by Elyce Winters. (Elyce also organized, together with the SBU Theatre Arts Department, a staged reading of Peter Parnell’s QED on April 26).

Part of the Committee’s task is to devise a new model for the Program administration, and then to recruit a new Director. Our current thoughts: the Director should be a half-time position, ideally occupied by a free-lance curator based in New York City, with interest and experience in the science/technology/art interface. Such people exist; recruiting one will require an increase in the Arts Program budget (at present, $25,000 annually is assigned to a Director’s stipend; this will have to go up by $20 to $25,000 in my estimation).

2. The SCG Arts Program Advisory Committee has been constituted with members of the Stony Brook campus and community who have experience with the interaction between art, science and technology:

Stephanie Dinkins, Art Department

George Hart, computer scientist and sculptor

John Lutterbie, Theatre Arts and Art Departments

Tony Phillips, Mathematics Department, Chair

Daniel Weymouth, Music Department

3. Exhibition plans have been blocked out for the rest of this year:

• (Now through June 2012) Mathematics in Stone and Bronze, sculptures by Heleman Ferguson. Curated by Peter Pantaleon. Formal opening will coincide with the installation of Ferguson’s 24-foot bronze Umbilic Torus.

• (July-August 2012) Oceans. Multimedia, curated by Flo Tarjan. This show may have an artist-in-residence component.

• (September-December 2012) Controlled Evidence, curated by Megs Schedel, SBU Dept. of Music.

4. Concerts. Weekly recitals are planned during the SCGP Graduate Summer School and the Summer Simons Workshop. Artists will include Leon Livshin, piano (solo and with other musicians) and Ray Anderson, trombone. Daniel Weymouth is committed to organizing a series of monthly student/faculty recitals during the academic year.

Respectfully submitted,

Tony Phillips, Arts Program Advisory Committee Chair